# Ceresa cycles of bielliptic Picard curves

### v1

@@ 3.1. Definitions
A bielliptic Picard curve $C$ over a field $k$ (of characteristic $\neq 2,3$) is a nice plane quartic curve over $k$ with affine model of the form
@@
[28, §2.2-3] Let $C_{a,b}$ and $C_{a^{\prime},b^{\prime}}$ be bielliptic Picard curves over $k$.
@@ 3.2. The Prym variety
[28, §2.5] $P$ is an abelian surface and $J$ is isogenous to $P\times E$.
3.3. Galois action on the Neron–Severi group
The main ingredient we need from [28] for our proof of Theorem 1.1 is an explicit description of the ${\rm Gal}_{k}$-action on the Neron–Severi group ${\rm NS}(P_{k^{\operatorname{sep}}})$, as well as the fact that $P_{k^{\operatorname{sep}}}$ has quaternionic multiplication.
@@ 3.3. Galois action on the Neron–Severi group
We need to pass to a cover of $S$ to "see" all endomorphisms of $\mathcal{P}$. Let $\tilde{S}\rightarrow S$ be the finite etale cover given by adjoining a primitive third root of unity $\omega$ and a sixth root of $\Delta_{a,b}$. More formally, consider the ring extension
$R={\rm\mathbb{Z}}[a,b,\Delta_{a,b}^{-1}]\rightarrow\tilde{R}={\rm\mathbb{Z}}[\omega][1/6,a,b,\Delta_{a,b}^{-1},\varepsilon]/(\varepsilon^{6}-\Delta_{a,b})$
and define $\tilde{S}\rightarrow S$ to be the induced morphism of schemes. This morphism is finite, etale, and Galois, with Galois group $G:={\rm Aut}(\tilde{S}/S)$ a dihedral group of order $12$. The group $G$ naturally acts on ${\rm End}^{0}(\mathcal{P})$. Let $\mathrm{triv}\colon G\rightarrow{\rm GL}_{1}({\rm\mathbb{Q}})$ be the trivial representation and let $\mathrm{std}\colon G\rightarrow{\rm GL}_{2}({\rm\mathbb{Q}})$ be a model for the reflection representation, i.e. the unique irreducible $2$-dimensional representation of $G$ defined over ${\rm\mathbb{Q}}$. See §1.7 for our notations concerning abelian schemes.
@@
The schemes $S$ and $\tilde{S}$ are integral and normal; let $K$ and $\widetilde{K}$ denote their function fields and let $P=\mathcal{P}_{K}$ denote the generic fiber of $\mathcal{P}$. By [17, §2, Lemma 1], the natural map ${\rm End}(\mathcal{P}_{\tilde{S}})\rightarrow{\rm End}(P_{\widetilde{K}})$ is an isomorphism. By [41, Tag 0BQM], the natural map $G={\rm Aut}(\tilde{S}/S)\rightarrow{\rm Gal}(\widetilde{K}/K)$ is also an isomorphism. It therefore suffices to prove all claims over the generic point ${\rm Spec}(K)$. By [28, Lemma 6.16], $P$ is geometrically simple. If $\bar{K}$ denotes an algebraic closure of $K$ then [28, Lemma 6.2] implies that ${\rm End}^{0}(P_{\bar{K}})$ is a discriminant $6$ quaternion algebra, and [28, Corollary 6.6] shows that ${\rm End}^{0}(P_{\bar{K}})={\rm End}^{0}(P_{\tilde{K}})$. The claim concerning ${\rm Hom}^{\operatorname{sym}}(\mathcal{P}_{\tilde{S}},\mathcal{P}^{\vee}_{\tilde{S}})\otimes{\rm\mathbb{Q}}={\rm Hom}(P_{\tilde{K}},P^{\vee}_{\tilde{K}})\otimes{\rm\mathbb{Q}}$ follows from [28, Corollary 6.8] and the fact that the latter is isomorphic to ${\rm NS}(P_{k^{\mathrm{sep}}})\otimes{\rm\mathbb{Q}}$. ∎
@@
In the decomposition ${\rm Hom}^{\operatorname{sym}}(\mathcal{P}_{\tilde{S}},\mathcal{P}^{\vee}_{\tilde{S}})\otimes{\rm\mathbb{Q}}=\mathrm{triv}\oplus\mathrm{std}$ of Theorem 3.4, $\mu_{3}$ acts trivially on $\mathrm{triv}$ and $\mu_{3}$ acts on std¯tensor-productnormal-$\mathrm{std}\otimes{\rm\overline{\mathbb{Q}}}$ as a direct sum of the two distinct nontrivial characters.
@@
The proof of [28, Corollary 6.8] shows that if we write $B=\text{span}\{1,i,j,ij\}$ with $i^{2}=-3$, $j^{2}=2$, $ij=-ji$ and $\omega=(-1+i)/2$, then ${\rm Hom}^{\operatorname{sym}}(\mathcal{P}_{\tilde{S}},\mathcal{P}^{\vee}_{\tilde{S}})\otimes{\rm\mathbb{Q}}\simeq\text{span}\{1,j,ij\}$, $\mathrm{triv}$ is spanned by $1$, $\mathrm{std}$ is spanned by $\{i,ij\}$ and $\omega\in\mu_{3}$ acts via conjugation by $\omega\in B$. The description of the $\mu_{3}$-action on $\mathrm{triv}$ and $\mathrm{std}$ now follow from an explicit calculation. ∎
@@ 4.1. Correspondences
We denote by $\mathsf{SmProj}(S)$ the category of smooth projective $S$-schemes. For $X$ in $\mathsf{SmProj}(S)$ the Chow groups ${\rm CH}^{p}(X)$ are well defined, see §1.7. Write ${\rm CH}^{p}(X;{\rm\mathbb{Q}}):={\rm CH}^{p}(X)\otimes_{{\rm\mathbb{Z}}}{\rm\mathbb{Q}}$ and if $X\rightarrow S$ has equidimensional fibers write $d(X/S)$ for the relative dimension of $X\rightarrow S$. If $X$ and $Y$ are in $\mathsf{SmProj}(S)$ we define the ${\rm\mathbb{Q}}$-vector space of correspondences by ${\rm Corr}(X,Y):={\rm CH}^{*}(X\times_{S}Y;{\rm\mathbb{Q}})$. We say a correspondence has degree $r$ if it lies in the subspace
@@
where $X=\sqcup_{i}X_{i}$ is a decomposition such that $X_{i}\rightarrow S$ is equidimensional. If $p\in{\rm Corr}^{r}(X,Y)$ and $q\in{\rm Corr}^{s}(Y,Z)$ are correspondences, denote by $q\circ p=\pi_{XZ,*}(\pi_{XY}^{*}p\cdot\pi_{YZ}^{*}q)\in{\rm Corr}^{r+s}(X,Z)$ their composition. If $f\colon X\rightarrow Y$ is an $S$-morphism write $\Gamma_{f}\subset X\times_{S}Y$ for its graph and ${}^{t}\Gamma_{f}\subset Y\times_{S}X$ for its transpose. Write $\Delta_{X/S}$ for the graph of the identity $X\rightarrow X$.
@@ 4.2. Relative Chow motives
We denote by $\mathsf{Mot}(S)$ the category of relative Chow motives with respect to graded correspondences over $S$, see [13, §1] for more details. The results of [13] are stated under the standing assumption that $S$ is smooth and quasi-projective over a field, but their results continue to hold in our more general set-up since intersection theory is developed in this generality, see [25, §1, Remark 1.1].
@@ 4.3. Base changing motives
Suppose that $T$ is another adequate scheme and $f\colon T\rightarrow S$ a morphism. There are two situations in which we have a good theory of pullbacks [^2] Pullbacks can be defined in greater generality, see for example [19, Example 20.1.2], but we restrict ourselves to the concrete cases we need in this paper. [19, §20.1]:
@@
    If $f$ is an lci morphism, there are Gysin homomorphisms $f^{*}\colon{\rm CH}^{p}(X)\rightarrow{\rm CH}^{p}(X_{T})$ for every $X/S$ in $\mathsf{SmProj}(S)$ compatible with flat pullbacks.
@@ 4.4. Descending direct summands of motives
Let $\pi\colon\tilde{S}\rightarrow S$ be a finite etale morphism of adequate schemes. Suppose additionally that it is Galois and write $G$ for the group of Deck transformations, so we have an isomorphism of $S$-schemes $g\colon\tilde{S}\rightarrow\tilde{S}$ for every $g\in G$. Using the pullback functor from §4.3, we obtain functors $g^{*}\colon\mathsf{Mot}(\tilde{S})\rightarrow\mathsf{Mot}(\tilde{S})$. Let $M,N$ be objects of $\mathsf{Mot}(S)$ and write $\tilde{M}=\pi^{*}M$, $\tilde{N}=\pi^{*}N$ for the corresponding objects of $\mathsf{Mot}(\tilde{S})$. Then we have canonical identifications $\tilde{M}=g^{*}\tilde{M}$ and $\tilde{N}=g^{*}\tilde{N}$ for all $g\in G$. It follows by functoriality that ${\rm Hom}(\tilde{M},\tilde{N})$ inherits a $G$-action.
@@
This is well known when $S={\rm Spec}(k)$ [39, Lemma 1.17]; we sketch the generalization of the classical proof to the relative setting. Write $M=(X/S,p,m)$ and $N=(Y/S,q,n)$. Then an element of ${\rm Hom}(\tilde{M},\tilde{N})^{G}$ is the same as a $G$-invariant element $\alpha$ of $q\circ{\rm CH}^{d(X/S)+m-n}(X\times_{S}Y;{\rm\mathbb{Q}})\circ p$. Choose a cycle $Z$ on $X_{\tilde{S}}\times_{\tilde{S}}Y_{\tilde{S}}$ representing $\alpha$. Then $\sum_{g\in G}g^{*}Z$ descends to a unique cycle $W$ of $X\times_{S}Y$. Moreover the rational equivalence class $[W]$ is independent of the choice of $Z$ representing $\alpha$. It follows that $\alpha\mapsto\frac{1}{|G|}[W]$ is a well-defined inverse to the base change map above. ∎
@@
Since direct sum decompositions correspond bijectively to a pair of orthogonal idempotents in the endomorphism ring, this immediately follows from Lemma 4.1. ∎
@@ 4.5. Group actions on motives
Let $M$ be a motive in $\mathsf{Mot}(S)$ and $G$ a finite group acting on $M$. Then every idempotent in the group algebra ${\rm\mathbb{Q}}[G]$ gives rise to an idempotent in $M$. For example, if $\rho$ is an irreducible (but not necessarily absolutely irreducible) representation defined over ${\rm\mathbb{Q}}$ and $e_{\rho}$ is the corresponding central idempotent, we get a direct summand $M_{\rho}$ of $M$, called the $\rho$-isotypic component. Applying this to $\rho=$ the trivial representation, we see that the motive of $G$-fixed points $M^{G}=M_{\rho}$ is well-defined and a direct summand of $M$.
@@ 4.6. Relative Artin motives
Assume that $S$ is connected. An Artin motive in $\mathsf{Mot}(S)$ is a direct summand of a motive of the form $\mathfrak{h}(X/S)$, where $X\rightarrow S$ in $\mathsf{SmProj}(S)$ has relative dimension zero or equivalently, is finite etale. When $S$ is the spectrum of a field, Artin motives are the same as Galois representations on finite-dimensional ${\rm\mathbb{Q}}$-vector spaces; we give a similar (presumably well known) description in the relative setting. Let $f\colon\tilde{S}\rightarrow S$ be a connected finite etale cover with Galois group $G={\rm Aut}(\tilde{S}/S)$. We say an Artin motive $M$ in $\mathsf{Mot}(S)$ is trivialized by $f$ if $f^{*}M\simeq\mathbf{1}_{\tilde{S}}^{\oplus n}$ for some $n\in{\rm\mathbb{Z}}_{\geq 0}$. In that case the ${\rm\mathbb{Q}}$-vector space ${\rm Hom}(\mathbf{1}_{\tilde{S}},f^{*}M)$ is $n$-dimensional (since $\tilde{S}$ is connected) and has a natural $G$-action, by the same reasoning as §4.4.
@@
Every Artin motive in $\mathsf{Mot}(S)$ is trivialized by some connected finite etale Galois cover $f\colon\tilde{S}\rightarrow S$. Fixing such an $f$, the assignment $M\mapsto V_{M}:={\rm Hom}(\mathbf{1}_{\tilde{S}},f^{*}M)$ induces an equivalence between the full subcategory of Artin motives in $\mathsf{Mot}(S)$ trivialized by $f$ and the category of ${\rm Aut}(\tilde{S}/S)$-representations on finite-dimensional ${\rm\mathbb{Q}}$-vector spaces.

### v2

@@ 3.1. Definitions
A bielliptic Picard curve $C$ over a field $k$ (of characteristic $\neq 2,3$) is a smooth projective plane quartic curve over $k$ with affine model of the form
@@
[32, §2.2-3] Let $C_{a,b}$ and $C_{a^{\prime},b^{\prime}}$ be bielliptic Picard curves over $k$.
@@ 3.2. The Prym variety
[32, §2.5] $P$ is an abelian surface and $J$ is isogenous to $P\times E$.
3.3. Galois action on the Néron–Severi group
The main ingredient needed from [32] for our proof of Theorem 1.1 is an explicit description of the ${\rm Gal}_{k}$-action on the Néron–Severi group ${\rm NS}(P_{k^{\operatorname{sep}}})$, as well as the fact that $P_{k^{\operatorname{sep}}}$ has quaternionic multiplication.
@@ 3.3. Galois action on the Néron–Severi group
We need to pass to a cover of $S$ to "see" all endomorphisms of $\mathcal{P}$. Let $\tilde{S}\rightarrow S$ be the finite étale cover given by adjoining a primitive third root of unity $\omega$ and a sixth root of $\Delta_{a,b}$. More formally, consider the ring extension
$R={\rm\mathbb{Z}}[1/6,a,b,\Delta_{a,b}^{-1}]\rightarrow\tilde{R}={\rm\mathbb{Z}}[\omega][1/6,a,b,\Delta_{a,b}^{-1},\varepsilon]/(\varepsilon^{6}-\Delta_{a,b})$
and define $\tilde{S}\rightarrow S$ to be the induced morphism of schemes. This morphism is finite, étale, and Galois, with Galois group $G:={\rm Aut}(\tilde{S}/S)$ a dihedral group of order $12$. The group $G$ naturally acts on ${\rm End}^{0}(\mathcal{P}_{\tilde{S}})$. Let $\mathrm{triv}\colon G\rightarrow{\rm GL}_{1}({\rm\mathbb{Q}})$ be the trivial representation and let $\mathrm{std}\colon G\rightarrow{\rm GL}_{2}({\rm\mathbb{Q}})$ be a model for the reflection representation, i.e. the unique irreducible $2$-dimensional representation of $G$ defined over ${\rm\mathbb{Q}}$. See §1.7 for our notations concerning abelian schemes.
@@
The schemes $S$ and $\tilde{S}$ are integral and normal; let $K$ and $\widetilde{K}$ denote their function fields and let $P=\mathcal{P}_{K}$ denote the generic fiber of $\mathcal{P}$. By [19, §2, Lemma 1], the natural map ${\rm End}(\mathcal{P}_{\tilde{S}})\rightarrow{\rm End}(P_{\widetilde{K}})$ is an isomorphism. By [46, Tag 0BQM], the natural map $G={\rm Aut}(\tilde{S}/S)\rightarrow{\rm Gal}(\widetilde{K}/K)$ is also an isomorphism. It therefore suffices to prove all claims over the generic point ${\rm Spec}(K)$. By [32, Lemma 6.16], $P$ is geometrically simple. If $\bar{K}$ denotes an algebraic closure of $K$ then [32, Lemma 6.2] implies that ${\rm End}^{0}(P_{\bar{K}})$ is a discriminant $6$ quaternion algebra, and [32, Corollary 6.6] shows that ${\rm End}^{0}(P_{\bar{K}})={\rm End}^{0}(P_{\tilde{K}})$. The claim concerning ${\rm Hom}^{\operatorname{sym}}(\mathcal{P}_{\tilde{S}},\mathcal{P}^{\vee}_{\tilde{S}})\otimes{\rm\mathbb{Q}}={\rm Hom}^{\operatorname{sym}}(P_{\tilde{K}},P^{\vee}_{\tilde{K}})\otimes{\rm\mathbb{Q}}$ follows from [32, Corollary 6.8] and the fact that the latter is isomorphic to ${\rm NS}(P_{k^{\mathrm{sep}}})\otimes{\rm\mathbb{Q}}$. ∎
@@
In the decomposition ${\rm Hom}^{\operatorname{sym}}(\mathcal{P}_{\tilde{S}},\mathcal{P}^{\vee}_{\tilde{S}})\otimes{\rm\mathbb{Q}}=\mathrm{triv}\oplus\mathrm{std}$ of Theorem 3.4, $\mu_{3}$ acts trivially on $\mathrm{triv}$ and $\mu_{3}$ acts on $\mathrm{std}\otimes{\rm\overline{\mathbb{Q}}}$ as a direct sum of the two distinct nontrivial characters.
@@
The proof of [32, Corollary 6.8] shows that if we write $B=\text{span}\{1,i,j,ij\}$ with $i^{2}=-3$, $j^{2}=2$, $ij=-ji$ and $\omega=(-1+i)/2$, then ${\rm Hom}^{\operatorname{sym}}(\mathcal{P}_{\tilde{S}},\mathcal{P}^{\vee}_{\tilde{S}})\otimes{\rm\mathbb{Q}}\simeq\text{span}\{1,j,ij\}$, $\mathrm{triv}$ is spanned by $1$, $\mathrm{std}$ is spanned by $\{i,ij\}$ and $\omega\in\mu_{3}$ acts via conjugation by $\omega\in B$. The description of the $\mu_{3}$-action on $\mathrm{triv}$ and $\mathrm{std}$ now follow from an explicit calculation. ∎
@@ 4.1. Correspondences
We denote by $\mathsf{SmProj}(S)$ the category of smooth projective $S$-schemes. For $X$ in $\mathsf{SmProj}(S)$ the Chow groups ${\rm CH}^{p}(X)$ are well defined, see §1.7. For the remainder of the paper write ${\rm CH}^{p}(X;{\rm\mathbb{Q}}):={\rm CH}^{p}(X)\otimes_{{\rm\mathbb{Z}}}{\rm\mathbb{Q}}$ and if $X\rightarrow S$ has equidimensional fibers write $d(X/S)$ for the relative dimension of $X\rightarrow S$. If $X$ and $Y$ are in $\mathsf{SmProj}(S)$ we define the ${\rm\mathbb{Q}}$-vector space of correspondences by ${\rm Corr}(X,Y):={\rm CH}^{*}(X\times_{S}Y;{\rm\mathbb{Q}})$. We say a correspondence has degree $r$ if it lies in the subspace
@@
where $X=\sqcup_{i}X_{i}$ is a decomposition such that $X_{i}\rightarrow S$ is equidimensional. If $p\in{\rm Corr}^{r}(X,Y)$ and $q\in{\rm Corr}^{s}(Y,Z)$ are correspondences, denote by $q\circ p=\pi_{XZ,*}(\pi_{XY}^{*}p\cdot\pi_{YZ}^{*}q)\in{\rm Corr}^{r+s}(X,Z)$ their composition. For an $S$-morphism $f\colon X\rightarrow Y$, let $\Gamma_{f}\subset X\times_{S}Y$ be its graph and let ${}^{t}\Gamma_{f}\subset Y\times_{S}X$ be the transpose. Write $\Delta_{X/S}$ for the graph of the identity $X\rightarrow X$.
@@ 4.2. Relative Chow motives
We denote by $\mathsf{Mot}(S)$ the category of relative Chow motives with respect to graded correspondences over $S$, see [15, §1] for more details. The results of [15] are stated under the standing assumption that $S$ is smooth and quasi-projective over a field, but their results continue to hold in our more general set-up since intersection theory is developed in this generality, see [29, §1, Remark 1.1].
@@ 4.3. Base changing motives
Suppose that $T$ is another adequate scheme and $f\colon T\rightarrow S$ a morphism. There are two situations in which we have a good theory of pullbacks [^1] Pullbacks can be defined in greater generality, see for example [21, Example 20.1.2], but we restrict ourselves to the concrete cases we need in this paper. [21, §20.1]:
@@
    If $f$ is a regular embedding, there are Gysin homomorphisms $f^{*}\colon{\rm CH}^{p}(X)\rightarrow{\rm CH}^{p}(X_{T})$ for every $X/S$ in $\mathsf{SmProj}(S)$ compatible with flat pullbacks.
@@ 4.4. Descending direct summands of motives
Let $\pi\colon\tilde{S}\rightarrow S$ be a finite étale morphism of adequate schemes. Suppose additionally that it is Galois and write $G$ for the group of Deck transformations, so we have an isomorphism of $S$-schemes $g\colon\tilde{S}\rightarrow\tilde{S}$ for every $g\in G$. Using the pullback functor from §4.3, we obtain functors $g^{*}\colon\mathsf{Mot}(\tilde{S})\rightarrow\mathsf{Mot}(\tilde{S})$. Let $M,N$ be objects of $\mathsf{Mot}(S)$ and write $\tilde{M}=\pi^{*}M$, $\tilde{N}=\pi^{*}N$ for the corresponding objects of $\mathsf{Mot}(\tilde{S})$. Then we have canonical identifications $\tilde{M}=g^{*}\tilde{M}$ and $\tilde{N}=g^{*}\tilde{N}$ for all $g\in G$. It follows by functoriality that ${\rm Hom}(\tilde{M},\tilde{N})$ inherits a $G$-action.
@@
This is well known when $S={\rm Spec}(k)$ [44, Lemma 1.17]; we sketch the generalization of the classical proof to the relative setting. Write $M=(X/S,p,m)$ and $N=(Y/S,q,n)$. Then an element of ${\rm Hom}(\tilde{M},\tilde{N})^{G}$ is the same as a $G$-invariant element $\alpha$ of $q\circ{\rm CH}^{d(X/S)+m-n}(X\times_{S}Y;{\rm\mathbb{Q}})\circ p$. Choose a cycle $Z$ on $X_{\tilde{S}}\times_{\tilde{S}}Y_{\tilde{S}}$ representing $\alpha$. Then $\sum_{g\in G}g^{*}Z$ descends to a unique cycle $W$ of $X\times_{S}Y$. Moreover the rational equivalence class $[W]$ is independent of the choice of $Z$ representing $\alpha$. It follows that $\alpha\mapsto\frac{1}{|G|}[W]$ is a well-defined inverse to the base change map above. ∎
@@
Since direct sum decompositions correspond bijectively to a pair of orthogonal idempotents in the endomorphism ring, this follows immediately from Lemma 4.1. ∎
@@ 4.5. Group actions on motives
Let $M$ be a motive in $\mathsf{Mot}(S)$ and $G$ a finite group acting on $M$. Then every idempotent in the group algebra ${\rm\mathbb{Q}}[G]$ gives rise to an idempotent in $M$. For example, if $\rho$ is an irreducible (but not necessarily absolutely irreducible) representation defined over ${\rm\mathbb{Q}}$ and $e_{\rho}$ is the corresponding central idempotent, we get a direct summand $M_{\rho}$ of $M$, called the $\rho$-isotypic component. Applying this to $\rho=$ the trivial representation, we have $e_{\rho}=\frac{1}{|G|}\sum_{g\in G}g$ and we see that the motive of $G$-fixed points $M^{G}=M_{\rho}$ is well-defined and a direct summand of $M$.
@@
4.6. Chow groups of motives
The (covariant) Chow groups of a motive $M$ in $\mathsf{Mot}(S)$ are by definition the ${\rm\mathbb{Q}}$-vector spaces ${\rm CH}^{p}(M):={\rm Hom}(\mathbb{L}^{\otimes p},M)$ for $p\in{\rm\mathbb{Z}}$, where we recall from §4.2 that $\mathbb{L}$ denotes the Lefchetz motive. For example, ${\rm CH}^{p}(\mathfrak{h}(X/S))={\rm CH}^{p}(X;{\rm\mathbb{Q}})$ for all $X/S$ in $\mathsf{SmProj}(S)$.
Suppose a finite group $G$ acts on $M$. This induces an action of $G$ on ${\rm CH}^{p}(M)$ for every $p$. If $e$ is an idempotent in ${\rm\mathbb{Q}}[G]$ corresponding to direct summand $e\cdot M$ of $M$, then ${\rm CH}^{p}(e\cdot M)=e\cdot{\rm CH}^{p}(M)$. In particular, ${\rm CH}^{p}(M^{G})={\rm CH}^{p}(M)^{G}$.
@@ 4.7. Relative Artin motives
Assume that $S$ is connected. An Artin motive in $\mathsf{Mot}(S)$ is a direct summand of a motive of the form $\mathfrak{h}(X/S)$, where $X\rightarrow S$ in $\mathsf{SmProj}(S)$ has relative dimension zero or equivalently, is finite étale. When $S$ is the spectrum of a field, Artin motives are the same as Galois representations on finite-dimensional ${\rm\mathbb{Q}}$-vector spaces; we give a similar (presumably well known) description in the relative setting. Let $f\colon\tilde{S}\rightarrow S$ be a connected finite étale cover with Galois group $G={\rm Aut}(\tilde{S}/S)$. We say an Artin motive $M$ in $\mathsf{Mot}(S)$ is trivialized by $f$ if $f^{*}M\simeq\mathbf{1}_{\tilde{S}}^{\oplus n}$ for some $n\in{\rm\mathbb{Z}}_{\geq 0}$. In that case the ${\rm\mathbb{Q}}$-vector space ${\rm Hom}(\mathbf{1}_{\tilde{S}},f^{*}M)$ is $n$-dimensional (since $\tilde{S}$ is connected) and has a natural $G$-action, by the same reasoning as §4.4.
@@
Every Artin motive in $\mathsf{Mot}(S)$ is trivialized by some connected finite étale Galois cover $f\colon\tilde{S}\rightarrow S$. Fixing such an $f$, the assignment $M\mapsto V_{M}:={\rm Hom}(\mathbf{1}_{\tilde{S}},f^{*}M)$ induces an equivalence between the full subcategory of Artin motives in $\mathsf{Mot}(S)$ trivialized by $f$ and the category of ${\rm Aut}(\tilde{S}/S)$-representations on finite-dimensional ${\rm\mathbb{Q}}$-vector spaces.